The temperature of a gas present in a closed vessel is increased by 5°C so that it`s pressure is increased by 1%.find out initial temperature of gas

Grade:12th pass

1 Answers

Arun
6 years ago
As vessel is closed , so volume of the container is constant.

P proportional to T
where P = pressure 
and T= temperature

so,
P1/P2 = T1/T2

now pressure in increased by 1%
Hence
P2 = P1 + P1*1/100 = 1.01 P1
T2 = T1+5

so,
P1/1.01P1 = T1/T1+5
T1+ 5 = 1.01 T1
0.01 T1 = 5
T1 = 500 K
T1 = 227°C